OpenStudy (anonymous):

There are five different types of bottles in a wine cellar. How many ways can four bottles be chosen from the cellar?

OpenStudy (anonymous):

120 WAYS

ganeshie8 (ganeshie8):

Number of ways of choosing 4 bottles = Number of ways of leaving 1 bottle behind

OpenStudy (anonymous):

5c4=5*4*3*2=120

ganeshie8 (ganeshie8):

you're using permutation formula but this question is about choosing bottles - so its a combination problem

OpenStudy (anonymous):

yes combination problem.

OpenStudy (anonymous):

it is confusing me.tell me how to do this simple problem

ganeshie8 (ganeshie8):

So we have 5 bottles, tell me how many ways you can leave one bottle behind ?

OpenStudy (anonymous):

p5=5! 4 bottles to be chosen=4*3!/5!

ganeshie8 (ganeshie8):

forget about formulas for few minutes, okay ?

ganeshie8 (ganeshie8):

see if you can answer my previous question

ganeshie8 (ganeshie8):